OpenStudy (anonymous):

simplify each radical expression. the radical to 75 + the radical to 3

OpenStudy (anonymous):

\[\sqrt{75}+\sqrt{3}\\ =\sqrt{25*3}+\sqrt{3}\\ =5\sqrt{3}+\sqrt{3}\\ =6\sqrt{3} \]